OpenStudy (anonymous):

evaluate lim x --> infinity of function (e ^ -x) ln x using l'hopitals rule Thanks to anyone who gives it a shot

OpenStudy (anonymous):

\[\LARGE \lim_{x\to \infty}e^{-x}\cdot \ln x\] is this what you mean ?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

\[\LARGE (u\cdot v)'=u'\cdot v+u\cdot v'\] you're supposed to know this rule... do you know it ?

OpenStudy (anonymous):

absolutely..but hopitals rule doesnt apply it..at least i didnt think it did

OpenStudy (anonymous):

...now you got me confused to.. O_O , is L'Hopital's Rule about derivatives when you're \[\LARGE \frac{\infty}{\infty} \quad \quad , \quad \quad \frac00\] ... ? :(

OpenStudy (experimentx):

0

OpenStudy (anonymous):

@experimentX yeah I know it's 0 .. http://www.wolframalpha.com/input/?i=lim_%7Bx%5Cto+%5Cinfty%7De%5E%28-x%29*ln+x but how to do it !! ??

OpenStudy (experimentx):

put that e^x on down and use L'hospital rue

OpenStudy (anonymous):

i got that far experiment..but then i get (1/x)/(e^x) after taking the derivative...at that point im stuck

OpenStudy (experimentx):

now .. as x-> inf your function = 1/inf = 0

OpenStudy (anonymous):

doesnt numerator go to 0 and denom to inf? (1/x --> 0 with x --> inf)

OpenStudy (experimentx):

yeah .. so it is your limit ...better than inf/inf isn't it??

OpenStudy (anonymous):

use drawing.. you can understand each other better !

OpenStudy (experimentx):

yeah .. just click on the link posted by @Kreshnik

OpenStudy (anonymous):

alright thanks alot..didnt see the link..take care

OpenStudy (experimentx):

oops sorry there wasn't graph!